How many grams of aluminum are needed to react with 2.08 kg of iron oxide?


1
Expert's answer
2022-03-27T10:30:05-0400
  • Write an equation: 2Al+3FeO---> Al2O3 + 3Fe
  • Convert 2.08kg into moles 72g =1mole, what about 2080g
  • 2080*1/72=28.8889moles
  • Find equivalent moles of Al by mole ratio
  • Al:FeO=2:3
  • If ratio 3=28.8889moles, what about ratio 2.
  • 2*28.8889/3=19.2593moles
  • Convert to mass
  • 27g = 1 mole, what about 19.2593moles
  • 19.2593*27 = 520g